### v2

# The Condorcet Dimension of Metric Spaces

Abstract
A Condorcet winning set is a set of candidates such that no other candidate is preferred by at least half the voters over all members of the set. The Condorcet dimension, which is the minimum cardinality of a Condorcet winning set, is known to be at most logarithmic in the number of candidates. We study the case of elections where voters and candidates are located in a $2$-dimensional space with preferences based upon proximity voting. Our main result is that the Condorcet dimension is at most $3$, under both the Manhattan norm and the infinity norm, natural measures in electoral systems.
1 Introduction
Consider an election with a set of candidates $\mathcal{C}$ of cardinality $m$ and a set of voters $\mathcal{V}$ of cardinality $n$, where each voter $v\in\mathcal{V}$ has a strict preference list $\succ_{v}$ over the candidates. An ideal winner in such an election is a Condorcet winner, a candidate that “beats” any other candidate in a pairwise voting contest. Formally, $i\in C$ is a Condorcet winner if, for every candidate $j\in\mathcal{C}\setminus\{i\}$, more than half of the voters prefer $i$ over $j$. To avoid ties, we assume the number $n$ of voters is odd. Unfortunately, it is easy to construct elections where a Condorcet winner does not exist; indeed typically there is no Condorcet winner.
Given this, Elkind, Lang and Saffidine [12] proposed a relaxation where, rather than a single winning candidate, we desire a set of winning candidates that collectively beats any other candidate. Specifically, a set of candidates $S\subseteq\mathcal{C}$ is a Condorcet winning set if, for every candidate $j\in\mathcal{C}\setminus S$, more than half of the voters prefer some (voter-dependent) candidate in $S$ to $j$. Elkind et al. [12] showed that a Condorcet winning set always exists provided we allow the winning set to be large enough.
Theorem 1 ([12]).
In any election, there is a Condorcet winning set of size at most $\lceil\log m\rceil$.
Proof.
Consider the majority digraph which contains a vertex for each candidate $i\in\mathcal{C}$ and an arc $(j,i)$ if a majority of voters prefer $i$ over $j$. Since the number of voters is odd, the majority digraph is a tournament. Megiddo and Vishkin [19] proved that any tournament contains a dominating set of size at most $\lceil\log m\rceil$. To see this, simply select the vertex with the largest number of in-neighbours; remove this vertex and its in-neighbours and recurse. A dominating set $T$ in the majority digraph is a Condorcet winning set. The theorem follows. (In fact, a dominating set $T$ has the stronger property that, for any candidate $j\in\mathcal{C}\setminus T$, there exists a specific candidate $i(j)\in T$ that more than half of the voters prefer over $j$. However, a Condorcet winning set does not need to be a dominating set in majority digraph.) ∎
Theorem 1 gives a logarithmic upper bound on the Condorcet dimension, the minimum cardinality of a Condorcet winning set. What about lower bounds? Since a Condorcet winner need not exist, there are elections with Condorcet dimension at least $2$. Further, Elkind et al. [12] presented instances with Condorcet dimension $3$; in addition, Geist [16] presented an instance with just six candidates and six voters of Condorcet dimension $3$. No instances with Condorcet dimension great than $3$ are currently known. However, [12] conjectured that elections exist with arbitrarily high Condorcet dimension.
1.1 The Spatial Model of Voting
This hypothesis by Elkind et al. [12] motivates our work. Specifically, we study Condorcet winning sets in the classical spatial model of voting where the voters and candidates are located in a metric space. We make the standard assumption of proximity voting, whereby each voter $v$ ranks the candidates based upon their distance to $v$, with closer candidates more preferred.
The spatial model of voting originates in the ground-breaking works of Hotelling [18] and Black [4], and was formalized by Downs [11]. These works all considered one-dimensional metric spaces. Of particular interest here is the median voter theorem, where Black [4] showed that the closest candidate to the median voter is a Condorcet winner. The study of higher dimensional metric spaces was instigated by Davis, Hinich, and Ordeshook [10]. The spatial model has since become ubiquitous in the study of voter and candidate behaviour, as illustrated by the books [13, 14, 20, 21, 22]. The spatial model has also recently attracted huge interest in computer science; see, for example, [3, 2, 7, 8, 15, 17, 23].
In spatial models of dimension greater than one, a choice of distance measure can be made. In this paper, we focus on the $p$-norms, [^1] For ${\bf z}\in\mathbb{R}^{D}$, the $p$-norm is defined as $||{\bf z}||_{p}=\sqrt[p]{z_{1}^{p}+z_{2}^{p}+\cdots+z_{D}^{p}}$. specifically the Manhattan norm ($p=1$) and the infinity norms ($p=\infty$). These two norms are most appropriate in political settings where each axis represents a distinct policy or characteristic. Under the Manhattan norm, a voter prefers the candidate whose sum of policy differences over all policies is minimized; that is, the voter desires a candidate close to them on average over all policies. Under the infinity norm, a voter prefers the candidate whose maximum difference over all policies is minimized; that is, the voter desires a candidate close to them on every policy.
It is known that any set of voter preferences over candidates has a corresponding embedding in a metric space with proximity voting, for the Euclidean norms [6] and Manhattan norms [9]. However, this embedding may require the dimension $D$ of the metric space to be high, specifically, $D=\min\,[m-1,n]$. We provide simple constructions that extends this conclusion to any $p$-norm, provided $D=\min\,[m,n]$.
This is in stark contrast to practice where the embedding is often in small dimension $D$; for example, [1] suggests that a small number of dimensions provide a good explanation for German electoral data. This observation motivates our study of low dimensional metric spaces, in particular $D=2$. Of course, if the dimension $D$ of the metric space is bounded then this does restrict the set of feasible preferences in the election. Indeed we show that this restriction guarantees that the Condorcet dimension is at most $3$ when $D=2$.
1.2 Our Results
We study metric spaces with bounded dimension $D$. Recall the median value theorem tell us that for $D=1$ a Condorcet winner always exists. What about the case $D=2$? In fact, Davis et al. [10] proved over half a century ago that a Condorcet winner need not exist in two dimensional metric spaces for the Euclidean norm. Consequently, it makes sense to study the Condorcet dimension for $D\geq 2$. Our main technical result, presented in Sections 2 and 3, is that for two dimensions the Condorcet dimension is at most $3$.
Meta-Theorem 1 (Theorem 3 and Theorem 5).
In a $2$-dimensional metric space (with Manhattan norm or infinity norm), the Condorcet dimension of any election is at most $3$.
Thus, the hypothesis of [12] does not hold when $D=2$. In Section 4, we further provide a lower bound of $2$, for any $p$-norm, by formalizing and instantiating the construction of [10].
{restatable*}
[]lemlb In a $2$-dimensional metric space under the $p$-norm, there exist instances with Condorcet dimension at least $2$, for any $p\geq 1$.
Finally, in Section 5 we extend the conclusions of [6] and [9] to any $p$-norm by showing that any set of voter preferences over candidates has a corresponding embedding in a metric space with proximity voting with $D=\min\,[m,n]$.
Meta-Theorem 2 (Theorem 6 and Theorem 7).
Given $m$ candidates and $n$ voters, an embedding of dimension $D=\min\,[m,n]$ can be computed in polynomial time, for any $p$-norm.
2 An Upper Bound on the Condorcet Dimension: Manhattan Norm
We begin with the Manhattan norm, whose distance function is the $1$-norm. Thus, given two points ${\bf p}_{1}=(x_{1},y_{1})$ and ${\bf p}_{2}=(x_{2},y_{2})$, the distance between them is $d^{1}({\bf p}_{1},{\bf p}_{2})=|x_{1}-y_{1}|+|x_{2}-y_{2}|$. Throughout this section, all notions of distance and closeness refer to the Manhattan norm; thus, we will omit the superscript and write $d({\bf p}_{1},{\bf p}_{2})=d^{1}({\bf p}_{1},{\bf p}_{2})$.
Let $\bar{x}$ be the median $x$-coordinate of the $n$ voters. Similarly, let $\bar{y}$ be the median $y$-coordinate of the voters. Without loss of generality, we may assume $\bar{x}=0$ and $\bar{y}=0$ by shifting all points accordingly.
The $x$-axis parallel line at $x=\bar{x}$ and the $y$-axis parallel line at $y=\bar{y}$ divide the plane into four quadrants $\{Q_{1},Q_{2},Q_{3},Q_{4}\}$. We label the quadrants in counter-clockwise order starting with the all-positive quadrant $Q_{1}$ (thus, $Q_{3}$ is the all-negative quadrant).
For each quadrant $Q_{i}$, $1\leq i\leq 4$, let $c_{i}$ be the candidate in $Q_{i}$ closest to the origin $(0,0)$. Note that $c_{i}$ will not exist if $Q_{i}$ is empty, but at least one of the $c_{i}$ exists as there are $m\geq 1$ candidates. Observe that the $c_{i}$ need not be distinct; indeed if $c^{*}=(0,0)$ is a candidate then $c_{1}=c_{2}=c_{3}=c_{4}=c^{*}$.
We claim that $S=\{c_{1},c_{2},c_{3},c_{4}\}$ is a Condorcet winning set.
Lemma 2.
If $c\in Q_{i}$ then at least half the voters prefer $c_{i}$ over $c$.
Proof.
Take any candidate $c=(x,y)$. Without loss of generality, let $c$ be in the positive quadrant $Q_{1}$. Since $c\in Q_{1}$, it is the case that $c_{1}=(x_{1},y_{1})\in Q_{1}$ exists. We may assume $c\neq c_{1}$. Since $c_{1}$ is the closest candidate to the origin in the positive quadrant $Q_{1}$, we have that $x+y\geq x_{1}+y_{1}$. Equivalently $x-x_{1}\geq y_{1}-y$. We now have two cases:
  1. 1.
    $y-x\leq y_{1}-x_{1}$
    We will show that at least half the voters prefer $c_{1}=(x_{1},y_{1})$ over $c=(x,y)$. Since $\bar{x}=0$ is the median $x$-coordinate, we know that at least half the voters have an $x$-coordinate at most zero. Take any such voter $v=(x_{v},y_{v})$ with $x_{v}\leq 0$, that is, any voter $v\in Q_{2}\cup Q_{3}$. We claim $d(v,c_{1})<d(v,c)$, that is, $v$ prefers $c_{1}$ over $c$. By the Manhattan metric, we have:
    $d(v,c_{1})$ $=$ $d(v,(0,y_{v}))+d((0,y_{v}),c_{1})$
    $=$ $|x_{v}|+d((0,y_{v}),c_{1})$
    $=$ $|x_{v}|+x_{1}+|y_{1}-y_{v}|.$
    Similarly,
    $d(v,c)$ $=$ $|x_{v}|+x+|y-y_{v}|.$
    Thus,
    $d(v,c)-d(v,c_{1})$ $=$ $x-x_{1}+|y-y_{v}|-|y_{1}-y_{v}|.$
    Recall $x-x_{1}\geq y_{1}-y$. Moreover, by the case assumption, we have $x-x_{1}\geq y-y_{1}$. Hence, $x-x_{1}$ is a positive number and thus, $x-x_{1}\geq|y-y_{1}|$. We conclude that
    $d(v,c)-d(v,c_{1})$$\geq$ $|y-y_{1}|+|y-y_{v}|-|y_{1}-y_{v}|$
    $\geq$ $0$
    The second inequality follows from the triangle inequality. Since the preferences orderings are strict, we have $d(v,c)-d(v,c_{1})>0$. Consequently, voter $v$ prefers candidate $c_{1}$ over candidate $c$, as desired.
  2. 2.
    $y-x\geq y_{1}-x_{1}$
    A similar argument applies with respect to voters of the form $v=(x_{v},y_{v})$ with $y_{v}\leq 0$, that is, any voter $v\in Q_{3}\cup Q_{4}$ prefers $c_{1}$ over $c$.
Given that each $c_{i}$, $1\leq i\leq 4$ beats every other candidate in its quadrant, we conclude that $S$ is indeed a Condorcet winning set. ∎
Let $S=\{c_{1},c_{2},c_{3},c_{4}\}$. We claim that three out of the four candidates already form a a Condorcet winning set.
Theorem 3.
In a $2$-dimensional metric space with Manhattan norm, the Condorcet dimension of any election is at most $3$.
Proof.
Let $S_{-i}=S\setminus\{c_{i}\}$ for each $1\leq i\leq 4$. If $S_{-i}$ is not a Condorcet winning set then there exists a candidate $b_{i}$ that more than half the voters prefer over the candidates in $S_{-i}$. Specifically, for more than half the voters, the closest candidate in $S_{-i}\cup\{b_{i}\}$ is $b_{i}$. We claim $b_{i}\in Q_{i}$. Assume $b_{i}\in Q_{j}$ with $j\neq i$, then by Lemma 2, at least half the voters prefer $c_{j}$ over $b_{i}$, a contradiction.
Next, consider a weighted directed graph $H$ on four vertices $\{1,2,3,4\}$, one for each of the four quadrants. Let the weight of vertex $i$ be $w_{i}$ be equal to the fraction of voters that rank $c_{i}$ highest in $S$. Hence, $w_{1}+w_{2}+w_{3}+w_{4}=1$. Let the weight of arc $(i,j)$, $i\neq j$ (i.e., no self-loops), be $w_{i,j}$ equal to the fraction of voters that prefer $c_{i}$ given the candidates in $S$, but prefer $b_{j}$ when given the candidates $S\cup\{b_{j}\}$. We omit the arc $(i,j)$ if $w_{i,j}=0$.
We claim that each vertex in $H$ has in-degree at most one. This follows from Lemma 2, as in both cases of the proof, every vertex in at least two of the three other quadrants prefer $c_{j}$ over $b_{j}$. This means, voters in those two quadrants cannot prefer $c_{i}$ in $S$ and prefer $b_{j}$ in $S\cup\{b_{j}\}$; otherwise they strictly prefer $b_{j}$ over $c_{j}$. Let $\ell(i)$ ($\neq i$ as there are no self-loops) be the quadrant with $w_{\ell(i),i}>0$ if it exists, or an arbitrary adjacent quadrant otherwise (This implies every vertex in $H$ has in-degree exactly one.).
Recall, we are assuming that $S_{-i}$ is not a Condorcet winning set and that $b_{i}$ is preferred by more that half the voters to $S_{-i}$. This implies that $w_{i}+w_{\ell(i),i}>0.5$, for all $1\leq i\leq 4$. This counts the fraction of voters that do prefer $b_{j}$ over candidate $c_{\ell(i)}$, and (potentially) overestimates that $b_{j}$ also gets all voters $c_{j}$ did receive previously.
Observe that $w_{\ell(i),i}\leq w_{\ell(i)}$ as the fraction of all voters preferring $b_{j}$ after its introduction over $c_{\ell(i)}$ cannot be more than the number that preferred $c_{\ell(i)}$ in the first place.
As $H$ has in-degrees exactly one, it contains a perfect matching (ignoring the arc directions). Without loss of generality, let this matching be $(2,1)$ and $(4,3)$. Then, we have
$w_{1}+w_{2}+w_{3}+w_{4}$$\geq$ $w_{1}+w_{2,1}+w_{3}+w_{4,3}$
$=$ $\left(w_{1}+w_{2,1}\right)+\left(w_{3}+w_{4,3}\right)$
$>$ $\frac{1}{2}+\frac{1}{2}~{}=~{}1~{}.$
This is a contradiction as it says that the overall weight of all candidates in $S$ is more than 1. This implies that at least one of the sets $S_{-i}$ is a Condorcet winning set of cardinality three. ∎
3 An Upper Bound on the Condorcet Dimension: Infinity Norm
Let us now consider the infinity norm (or supremum norm). Given two points, ${\bf p}_{1}=(x_{1},y_{1})$ and ${\bf p}_{2}=(x_{2},y_{2})$ in two dimensions, the distance between them is $d^{\infty}({\bf p}_{1},{\bf p}_{2})=\max(|x_{1}-x_{2}|,|y_{1}-y_{2}|)$. The Manhattan norm and the infinity norm are closely related. Specifically, imagine that we rotate the axes of measurement by $45$ degrees. Let $\hat{d}^{\infty}$ be the infinity norm using these new axes of measurement. The resultant distances are equivalent to the Manhattan norm.
Claim 4.
For any pair of points ${\bf p}_{1}$ and ${\bf p}_{2}$ in two dimensions, we have $d^{1}({\bf p}_{1},{\bf p}_{2})=\sqrt{2}\cdot\hat{d}^{\infty}({\bf p}_{1},{\bf p}_{2})$.
Proof.
Without loss of generality, we may assume that ${\bf p}_{1}={\bf 0}$ and that ${\bf p}_{2}$ lies in the positive quadrant. Thus $d^{1}({\bf 0},{\bf p}_{2})=|x_{2}|+|y_{2}|=x_{2}+y_{2}$. On the other hand, $\hat{d}^{\infty}({\bf 0},{\bf p}_{2})=\frac{1}{\sqrt{2}}\cdot(x_{2}+y_{2})$. The claim follows. ∎
Theorem 5.
In a $2$-dimensional metric space with infinity norm, the Condorcet dimension of any election is at most $3$.
Proof.
Take any election instance. By Claim 4, there is a equivalent election with the Manhattan norm that induces identical preference lists. By Theorem 3, this election has Condorcet dimension at most $3$. ∎
4 A Lower Bound on the Condorcet Dimension
In this section, we study lowers bounds on the Condorcet dimensions. Again, we focus on the case $D=2$ of metric spaces in two dimensions. Specifically, we prove that instances exist in two dimensions where a Condorcet winner does not exist. This result applies for any $p$-norm, including the infinity and Manhattan norms.
[…]
Proof.
Consider an election with three voters ${\bf v}_{1}=(9,0),{\bf v}_{2}=(0,9)$ and ${\bf v}_{3}=(-9,0)$ and three candidates ${\bf p}_{1}=(1,-1),{\bf p}_{2}=(8,10)$ and ${\bf p}_{3}=(-9,9)$. Using the $p$-norm, for any $p\geq 1$, the preferences rankings of the voters are ${\bf v}_{1}:{\bf p}_{1}\succ{\bf p}_{2}\succ{\bf p}_{3}$, ${\bf v}_{2}:{\bf p}_{2}\succ{\bf p}_{3}\succ{\bf p}_{1}$, and ${\bf v}_{3}:{\bf p}_{3}\succ{\bf p}_{1}\succ{\bf p}_{2}$. Hence there is a Condorcet cycle and so no Condorcet winner. ∎
5 Embedding Voter Preferences in a High Dimensional Metric Space
Bogomolnaia and Laslier [6] and Chen at al.[9] showed any set of voter preferences over candidates has a corresponding embedding in a metric space with proximity voting, for the Euclidean norm ($p=2$) and Manhattan norm ($p=1$), respectively, for $D=\min\,[m-1,n]$. We provide simple constructions for general $p$-norms, giving $D=\min[m,n]$, in the subsequent two theorems.
Theorem 6.
Given $m$ candidates and $n$ voters, each with a strict preference list $\succ_{v}$ per voter $v$ over the candidates. An embedding of dimension $D=m$ can be computed in polynomial time for any $p$-norm, $p\geq 0$.
Proof.
In our construction we place voter $j$ at position ${\bf y}_{j}$ and candidate $i$ at position ${\bf x}_{i}$. First, we place the $m$ candidates at the corners of a simplex in $m$ dimensions. Specifically, let ${\bf e}_{i}=(0,\dots,0,1,0,\dots,0)$, where the $i$th coordinate has value $1$. Then we position candidate $i$ at ${\bf x}_{i}=2m\cdot{\bf e}_{i}$, for each $1\leq i\leq m$. Next, consider voter $j$ for all $1\leq j\leq n$. Let $\rho(i,j)$ be the position of candidate $i$ in the preference list of voter $j$. Place voter $j$ at ${\bf y}_{j}=(m-\rho(1,j),m-\rho(2,j),\dots,m-\rho(m,j))$. Clearly this embedding can be computed in polynomial time.
Now, for the infinity norm we have
$d^{\infty}({\bf y}_{j},{\bf x}_{i})$ $=$ $\max[(2m-(m-\rho(i,j)),\max_{\ell\neq i}(m-\rho(\ell,j)]$
$=$ $\max[m+\rho(i,j)),\max_{\ell\neq i}(m-\rho(\ell,j)]$
$=$ $m+\rho(i,j)$
Observe that these distances are increasing with $\rho(i,j)$. Hence this positioning is consistent with the preference ranking of voter $j$.
Moreover this construction also works for any $p$-norm, with $0\leq p<\infty$. To see this, observe that
$d^{p}({\bf y}_{j},{\bf x}_{i})^{p}-d^{p}({\bf y}_{j},{\bf x}_{k})^{p}$ $=$ $(m+\rho(i,j))^{p}+\sum_{\ell\neq i}\big{(}m-\rho(\ell,j)\big{)}^{p}-(m+\rho(k,j))^{p}-\sum_{\ell\neq k}\big{(}m-\rho(\ell,j)\big{)}^{p}$
$=$ $\Big{(}(m+\rho(i,j))^{p}-(m+\rho(k,j)^{p}\Big{)}+\Big{(}\big{(}m-\rho(k,j)\big{)}^{p}-\big{(}m-\rho(i,j)\big{)}^{p}\Big{)}$
$>$ 00 0
where the strict inequality holds whenever $\rho(i,j)>\rho(k,j)$. Thus the distances increase with $\rho(i,j)$. ∎
Theorem 7.
Given $m$ candidates and $n$ voters, each with a strict preference list $\succ_{v}$ per voter $v$ over the candidates. An embedding of dimension $D=n$ can be computed in polynomial time for any $p$-norm, $p>1$.
Proof.
In this construction, rather than the candidates, we place the $n$ voters at $n$ corners of an $n$-dimensional simplex. Specifically, we may use the construction of Bogomolnaia and Laslier [6]. Place voter $j$ at ${\bf y}_{j}=B\cdot{\bf e}_{j}$, for each $1\leq j\leq n$, where $B$ (to be determined) is large. Next we position the $m$ candidates. Again, let $\rho(i,j)$ be the position of candidate $i$ in the preference list of voter $j$. Then we position each candidate $i$ at ${\bf x}_{i}=(-\rho(i,1),-\rho(i,2),\dots,-\rho(i,n))$.
For the infinity norm, we then have
$d^{\infty}({\bf y}_{j},{\bf x}_{i})$ $=$ $\max\,\Big{[}\,B+\rho(i,j)\,,\,\max_{\ell\neq j}\rho(i,\ell)\big{)}\,\Big{]}$
$=$ $B+\rho(i,j)$
Here the second equality holds if $B\geq m$. These distances are then increasing with $\rho(i,j)$, so they are consistent with the preference ranking of voter $j$.
Next consider any $p$-norm, with $1<p<\infty$. Then
$d^{p}({\bf y}_{j},{\bf x}_{i})^{p}-d^{p}({\bf y}_{j},{\bf x}_{k})^{p}$ $=$ $(B+\rho(i,j))^{p}+\sum_{\ell\neq j}\rho(i,\ell)^{p}-(B+\rho(k,j))^{p}-\sum_{\ell\neq j}\rho(k,\ell)^{p}$
$\leq$ $(B+\rho(i,j))^{p}-(B+\rho(k,j))^{p}+(n-1)\cdot(M^{p}-1)$
$<$ 00 0
Here, since $p>1$, there exists a sufficiently large choice of $B$ such that the strict inequality holds whenever $\rho(i,j)<\rho(k,j)$. Ergo, the distances are increasing with $\rho(i,j)$ and are consistent with the preference ranking of voter $j$. ∎
6 Conclusion
We presented bounds on the Condorcet dimension under the spatial model of voting. Several open problems remain. For the case ${D=2}$, we have shown the Condorcet dimension is at least $2$ and at most $3$. Do any instances exist with Condorcet dimension $3$, or is the Condorcet dimension at most $2$ for $2$-dimensional metric spaces? Our upper bound holds for the Manhattan and infinity norms; does it extend to other distance norms? However, the outstanding open problem concerns the general case (equivalently, instances where $D$ can be arbitrarily large). Does the hypothesis of Elkind et al. [12], that the Condorcet dimension can be arbitrarily large, hold? We predict not. Specifically, we conjecture the Condorcet dimension is at most three in any election. The current evidence for this is an observation by Bloks [5] that elections where the majority digraph defines a tournament with minimum dominating set of size four seem to be “random” enough to have small Condorcet dimension.
Acknowledgement. We are very grateful to Jannik Peters for comments and for directing us to the papers [6] and [9].